OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (-2, 1). Which equation defines circle P?

OpenStudy (anonymous):

A. (x + 2)2 + (y – 1)2 = 8 B. (x – 2)2 + (y + 1)2 = 8 C. (x + 2)2 + (y – 1)2 = 64 D. (x – 2)2 + (y + 1)2 = 64

OpenStudy (anonymous):

the equation of a circle with center (a,b) and radius r is: (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

I'm really confused and I don't know how to use the formula. Could you explain?

OpenStudy (agent47):

(x-(-2))^2+(y-1)^2=8^2

OpenStudy (agent47):

That's what he/she means.

OpenStudy (anonymous):

\[(x-a)^{2}+(y-b)^{2}=r^{2}\]

OpenStudy (anonymous):

where (a,b) is the center of your circle and r is the radius

OpenStudy (anonymous):

In your case a=-2, b=1, and r = 8 put these into the general equation of a circle that I've given you

OpenStudy (anonymous):

How would I find x and y? I'm not good at equations like this "/

OpenStudy (agent47):

You don't need x and y... you just need to see where in your equation are a, b, and r

OpenStudy (agent47):

I don't even think you're trying to understand what we're saying.. Come on, pay attention..

OpenStudy (anonymous):

okay, thanks. I get it now.

OpenStudy (anonymous):

It's not as hard as I thought

OpenStudy (agent47):

An equation of a circle is: (x-h)^2+(y-k)^2=r^2 Where (h;k) is the center and r is the radius. Your center is (-2;1), radius is 8 Plug everything back into the equation of a circle to get the result.

OpenStudy (anonymous):

Thanks Agent47